OpenStudy (anonymous):

I need help on this Math Quiz:) Can someone help me? Only 7 questions:) PLEASE HELP!!!!

OpenStudy (anonymous):

ask them...

OpenStudy (anonymous):

Ask who?

OpenStudy (anonymous):

ask here... :)

OpenStudy (anonymous):

Oh lol So you can help me:) Okay the questions are: 1. Which expression is equivalent to (-2)(a+6)?

OpenStudy (anonymous):

-2a - 12

OpenStudy (anonymous):

(-2xa) + (-2 x 6)

OpenStudy (anonymous):

Thanks:) Next.... 2. To which subsets of real numbers does the number -22 belong?

OpenStudy (anonymous):

Can you help?

OpenStudy (anonymous):

Z

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i meant Set of integers

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

3. What is an algebraic expression for the following word phrase: "6 time the b and p?

OpenStudy (anonymous):

THA

OpenStudy (anonymous):

@AkashdeepDeb please come...He can help you because I can't understand some of questions...

OpenStudy (akashdeepdeb):

Tell me the question

OpenStudy (anonymous):

@Megan1999

OpenStudy (anonymous):

5. Evaluate u + xy, if u=18, x=10, and y=8.

OpenStudy (anonymous):

18 + (10x8)

OpenStudy (anonymous):

18 + 80 = 98

OpenStudy (anonymous):

Thank you! This is an essay question if you dont mind.... A square field has an area of 479ft^2. What is the approximate length of a side of the field? Give your answer your answer to the nearest foot. Explain your response.

OpenStudy (anonymous):

\[\sqrt{479^{2}} = 479\]

OpenStudy (anonymous):

Anyone??

OpenStudy (anonymous):

The next !

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."